OpenStudy (anonymous):

Amy is completing a craft project that involves covering only the lateral surface of a cylindrical container with fabric. The cylinder has a height of 11.1 in. and a radius of 5.9 in. To the nearest square unit, how much fabric does she need for this project?

OpenStudy (anonymous):

Fabric only needs to cover lateral side, not to upper or lower surface. So you just have to calculate the circumference of circle-shaped bottom and "lift" it up by height \[2r \pi * h = 2*5.9*\pi*11.1 = 411.48 \] The answer is 411.48 in^2.